Solve: Y” – 134 + 654 – 125y = 0 y(0) = 3, y’(0) = 19, y(0) = 87 g(t) = Preview
Answer #1

y-134 + 65 y-125y = 0 m²_13 m² +65m - 125=0 а) і че да уң mas, 4+ 31, y(t) = ciest tout lelax (32) + casin (34)). yo): 320